OpenStudy (anonymous):

Which set of side lengths will not form a right triangle? 5, 12, 13 7, 24, 25 9, 16, 20 8, 15, 17

OpenStudy (anonymous):

For this sum of squares of two sides will be equal to square of the largest side.. So check for third one: is 9^2 + 16^2 = 20^2 ???

OpenStudy (anonymous):

16+9 is 25

OpenStudy (anonymous):

I am saying to first square them: 9^2 + 16^2 = 20^2 ???

OpenStudy (anonymous):

its 400

OpenStudy (anonymous):

19 and 6 is337

OpenStudy (anonymous):

So it is not equal so this is not the right triangle..

OpenStudy (anonymous):

oooooo then its a